OpenStudy (anonymous):

H-E-L-P : HOW TO FIND DERIVATIVE SQUARE ROOT OF 8-6*X^2

OpenStudy (anonymous):

\[\sqrt{8-6*X^2}\]

OpenStudy (anonymous):

the derivative of \(\sqrt{f(x)}\) is \[\frac{f'(x)}{2\sqrt{f(x)}}\]

OpenStudy (anonymous):

can we do this step by step please.

OpenStudy (anonymous):

the steps are what i wrote above take the derivative of \(8-6x^2\) and by the power rule, you get \(-12x\) that goes in the numerator in the denominator you put \(2\sqrt{8-6x^2}\)

OpenStudy (anonymous):

oh... why do we put that as the denominator?

OpenStudy (anonymous):

all that is left to do is cancel a 2 top and bottom

OpenStudy (anonymous):

the derivative of \(\sqrt{x}\) is \(\frac{1}{2\sqrt{x}}\) therefore by the chain rule the derivative of \(\frac{1}{f}\) is \(\frac{f'}{2\sqrt{f}}\) where \(f\) is any function

OpenStudy (anonymous):

for example, the derivative of \(\sqrt{\sin(x)}\) is \[\frac{\cos(x)}{2\sqrt{\sin(x)}}\]

OpenStudy (anonymous):

the derivative of \[\sqrt{3x^2+2x}\] is \[\frac{6x+2}{2\sqrt{3x^2+2x}}\]

OpenStudy (anonymous):

after i have -12x/ 2radical 8-6x^2.. how can i find the critical points?